SpaceRef

SpaceRef


Hubble Space Telescope Advanced Camera for Surveys (ACS) Status Report #2

Status Report From: Space Telescope Science Institute
Posted: Friday, October 13, 2006

image

On Wednesday, October 4, the Advanced Camera for Surveys (ACS) Anomaly Review Board approved an on-orbit test to attempt recovery of the full operational capabilities of the High Resolution Channel (HRC).

The proposed on-orbit test involved toggling the HRC Analog Signal Processing Card 2 relay that was thought to be the most likely location of the open in the +35V bias line that caused the instrument to suspend operations on September 23.

Early on October 9 engineers sent commands to Hubble Space Telescope to toggle the suspect relay. Telemetry confirmed that the relay cycled open and closed as expected. Engineers determined this action succeeded in restoring the HRC at 5:40 am on October 9 during the first opportunity to restore power to the +35V bias line. Since this test was fully successful, further workarounds will not be required.

NASA engineers believe the cause of the open circuit was a tiny particle of dust or fabric physically interfering with the electrical contact.

Engineers will monitor performance of the HRC for the remainder of the week. If its performance remains stable HRC science observations will resume the week of October 15.

Additional status reports will not be issued if science observations begin next week as planned.

NASA Hubble Space Telescope Daily Report #4215

SIGNIFICANT EVENTS:

Flash Report for ACS HRC and CEB Single Toggle test

At 282/21:55 the ACS HRC CEB internal ASPC2 relay was successfully cycled from side 2 to side 1 and then back via real-time command (OPS Request #17938). All other HRC CEB internal relays were also successfully cycled.

ACS will start its transition from it monthly anneal cycle to WFHROper at 283/09:01. LVPS CEB power will be applied to both the HRC and WFC CEB at 283/09:40:33. At this time, the results of the single toggle will be seen in the telemetry down link and another flash report will be sent out indicating Joy/NoJoy.

Single Toggle test Flash Report #2

The ACS HRC CEB ASPC2 +35 voltage has been restored as a result of the single toggle. All other voltages and temperatures are within their normal ranges. The SISE will continue to monitor the input voltage over the next few days.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks